# Anderson localization for the multi-frequency quasi-periodic CMV matrices and quantum walks

### v1

@@ 3.3 Estimate for the Lyapunov exponent
Assume $\omega\in\mathbb{T}^{d}(p,q)$, $z\in\partial\mathbb{D}$, and $L(\omega,z)>\gamma>0$. Then for any $n\geq 2$,
@@
Obviously, $0\leq L_{n}(\omega,z)\leq C(z)$ for all $n$. Let $k$ be a positive integer such that $k\gamma>16C(z)$. Given $n>10$, let $l=[C_{1}(\log n)^{1/\sigma}]$ with some large $C_{1}$. Consider the integers $l_{0},2l_{0},\ldots,2^{k}l_{0}$. Then there exists some $0\leq j<k$ such that (with $l_{j}=2^{j}l_{0}$)
@@
It is not difficult to check that $M(\omega,z;x)$ is conjugate to an $SL(2,\mathbb{R})$ matrix $T(\omega,z;x)$ through the matrix
$Q=-\frac{1}{1+i}\left[\begin{array}[]{cc}1&-i\\ 1&i\end{array}\right]\in\mathbb{U}(2);$
@@
$T(\omega,z;x)=Q^{*}M(\omega,z;x)Q\in SL(2,\mathbb{R}),$
@@
Now we apply the avalanche principle (AP) to the matrices $B_{j}=T_{l}(\omega,z;x+(j-1)l\omega)$ for $1\leq j\leq m=[n/l]$ with $\mu=\exp(l\gamma/2)$. Note that $\mu>n^{2}$ if $C_{1}>4/\gamma$. According to (LDT), the phases $x$ such that
@@
Based on (AP), we have that
@@
We can finally prove this lemma by summing over $2^{k}n$ ∎
@@ 3.4 Green’s function and Poisson’s formula
$\Theta_{n}=\left[\begin{array}[]{cc}\overline{\alpha}_{n}&\rho_{n}\\ \rho_{n}&-\alpha_{n}\end{array}\right].$
@@
$\mathcal{L}_{+}=\left[\begin{matrix}\Theta_{0}&~{}&~{}\\ ~{}&\Theta_{2}&~{}\\ ~{}&~{}&\ddots\end{matrix}\right],\quad\mathcal{M}_{+}=\left[\begin{matrix}\mathbf{1}&~{}&~{}\\ ~{}&\Theta_{1}&~{}\\ ~{}&~{}&\ddots\end{matrix}\right].$
@@
We let $\mathcal{E}_{[a,b]}$ denote the restriction of an extended CMV matrix to a finite interval $[a,b]\subset\mathbb{Z}$, defined by
@@
For $z\in\mathbb{C}$, $\beta,\eta\in\partial\mathbb{D}$, we can define the characteristic determinant of matrix [$\mathcal{E}^{\beta,\eta}_{[a,b]}$,
@@
Since the equation ~$\tilde{\mathcal{E}}u=zu$ is equivalent to $(z\mathcal{L}^{*}-\mathcal{M})u=0$, the associated finite-volume Green’s functions are as follows:
@@
According to [18, Proposition 3.8] and [29, Section B.1], for $\beta,\eta\in\partial\mathbb{D}$, the Green’s function has the expression:
@@
From [18, Lemma 3.9], if $u$ satisfies ~$\tilde{\mathcal{E}}u=zu$, Poisson’s formula reads
@@
Without loss of generality, we restrict the extended CMV matrix to the interval $[0,n-1]\subset\mathbb{Z}$. Let $\varphi^{\beta,\eta}_{[0,n-1]}(\omega,z;x)=\det\big{(}z-\mathcal{E}^{\beta,\eta}_{[0,n-1]}\big{)}$ be the characteristic determinant of matrix [$\mathcal{E}^{\beta,\eta}_{[0,n-1]}$. It follows from [24, Theorem 2] that the relation between the $n$-step transfer matrix and the characteristic determinant is
@@
Based upon the above analysis, we can get the Green’s function estimate which will be of great significance in the proof of the Anderson localization.
@@
where $C_{0}$ and $\sigma$ are as in (LDT). Furthermore, for any $x$ outside $\Omega$,
@@
Without loss of generality, we assume that $10\leq j\leq k\leq n-$ 1 and $n\gg N_{0}(\gamma)$. Then one can obtain that
@@ 3.5 Semialgebraic sets
4 Proof of the main result
For easier reading, we recall the paving property [25] which we will use in the proof of the main result.
For $\Lambda\subset\mathbb{Z}$, we let
$\mathcal{E}_{\Lambda}=R_{\Lambda}\mathcal{E}R_{\Lambda}^{*},$
where $R_{\Lambda}$ is the restriction operator and
$G_{\Lambda}=(z-\mathcal{E}_{\Lambda})^{-1}.$
Recall the resolvent identity,
$G_{\Lambda}=(G_{\Lambda_{1}}+G_{\Lambda_{2}})-(G_{\Lambda_{1}}+G_{\Lambda_{2}})(\mathcal{E}_{\Lambda}-\mathcal{E}_{\Lambda_{1}}-\mathcal{E}_{\Lambda_{2}})G_{\Lambda},$
where $\Lambda$ is a disjoint union $\Lambda=\Lambda_{1}\cup\Lambda_{2}$ provided that the inverses make sense. One of the consequences of the resolvent identity is the following paving property.
We let $I\subset\mathbb{Z}$ be an interval of size $N>n$ such that for each $x\in I$, there is an interval $I^{\prime}\subset I$ of size $n$ satisfying
$\big{\{}y\in I:|x-y|<\frac{n}{10}\big{\}}\subset I^{\prime}$
and
$|G_{I^{\prime}}(n_{1},n_{2})|<e^{-c|n_{1}-n_{2}|}$
for some constant $c>0$ and $n$ sufficiently large. Then we also have that
$|G_{I}(n_{1},n_{2})|<e^{-\frac{c}{2}|n_{1}-n_{2}|}.$
For the proof of this statement, one can consult part (IV) of [2].
Proof of Theorem 1.1 Fix $n_{0}$ and consider the property
@@ 4 Proof of the main result
for all $10\leq m_{1},m_{2}\leq n_{0}-$ 1. Replace (4.1) by the condition
$\mathop{\sum}\limits_{0\leq m_{1},m_{2}\leq n_{0}-1}$ $e^{2\gamma|m_{1}-m_{2}|}\big{|}\mathrm{det}[(m_{1},m_{2})-\mathrm{minor}\,\,\mathrm{of}\,\,(z-\mathcal{E}_{[0,n_{0}-1]}^{\beta,\eta}(x,\omega))]\big{|}^{2}$
@@
For the condition (4.1), its left-hand side is a polynomial in ($\mathrm{Re}\alpha(x),\mathrm{Im}\alpha(x),\mathrm{Re}z,\mathrm{Im}z$). Since $\mathrm{Re}\alpha(x)$ and $\mathrm{Im}\alpha(x)$ are real analytic functions, then we can replace them by the trigonometric polynomials
@@
Fix $\omega\in\mathbb{T}^{d}(p,q)$ and $z$. According to Lemma 3.5, the exceptional set $\Omega$ does not only satisfy
@@
Fix $x_{0}\in\mathbb{T}^{d}$. Now we consider the orbit $\{x_{0}+j\omega:|j|\leq n_{1}\}$. In this case, we let $n_{1}=n_{0}^{C}$, where $C$ is a sufficiently large constant. Apply Lemma 3.7 with $\mathcal{S}=\Omega$, $B=C_{2}n_{0}^{3}$, $\eta=\exp(-C_{0}n^{\sigma})$ and $n=n_{1}$. Then the statement (3.15) implies that except for at most $n_{1}^{1-\delta}$ values of $|j|<n_{1}$, taking $x=x_{0}+j\omega$, one of the intervals $\Lambda$ from (4.6) satisfies (4.7).
Assume that $u=\{u(n)\}_{n\in\mathbb{Z}}$ satisfies the equation ~$\tilde{\mathcal{E}}(x_{0},\omega)u=zu$ with $u(0)=1$ and $|u(n)|\lesssim n^{C}$.
@@
If $-j_{0}+1$, $-j_{0}+2$, $j_{0}-2$, $j_{0}-1$ satisfy (4.8), one can obtain that
@@
Therefore, fixing $z$, if there is a state $u$ with $u(0)=1$, then for any large $n_{0}$, there exists some $j_{0}$, $|j_{0}|<n_{1}=n_{0}^{C}$ for which (4.10) holds.
Denote $\Sigma_{\omega}=\mathop{\cup}\limits_{|j|\leq n_{1}}\sigma(\mathcal{E}_{-j,j}^{\beta,\eta}(x_{0}))$. It follows from (4.7) and (4.10) that if
@@
Now consider the interval $[\frac{n_{2}}{2},2n_{2}]$ with $n_{2}=n_{0}^{C^{\prime}}$ ($C^{\prime}$ a sufficiently large constant). Suppose that we ensured that
@@
Invoking the paving property, one can obtain that
@@
Now we verify the assumption (4.13). For $|j|\leq n_{1}$, we consider the set
=$\mathcal{B}=\{(\omega,z^{\prime},x):\omega\in\mathbb{T}^{d}(p,q),z^{\prime}\in\sigma(\mathcal{E}_{[-j,j]}^{\beta,\eta}(x_{0})),x\in\Omega(z^{\prime})\}\subset\mathbb{T}^{d}\times\partial\mathbb{D}\times\mathbb{T}^{d}.$
@@
$[0,1]^{d}=\mathop{\bigcup}\limits_{l}\Big{(}x_{l}+[0,\frac{1}{n_{2}}]^{d}\Big{)},\,\,l\leq n_{2}^{d}.$
@@
Summing these sets over $j$, $j\leq n_{1}$, we get an $\omega$-set of measure at most $n_{1}n_{2}^{-\frac{1}{12}}<n_{2}^{-\frac{1}{13}}$. This implies that (4.13) holds for $\omega$ outside the $\omega$-set we obtained. Since we initially fixed $n_{0}$ and set $n_{1}=n_{0}^{C}$, $n_{2}=n_{0}^{C^{\prime}}$, we can denote the above $\omega$-set to be $\Omega_{n_{0}}$. Then we have that
@@
Let $\Omega_{*}=\mathop{\cup}\limits_{n}\mathop{\cap}\limits_{n_{0}>n}\Omega_{n_{0}}$. It follows that $\mathrm{mes}\Omega_{*}=0$ and Anderson localization holds for $\omega\in\mathbb{T}^{d}(p,q)\backslash\Omega_{*}$.
@@
5 Quantum walks
Quantum walks share a structural similarity with CMV matrices via unitary equivalence. Now we recall some basic knowledge about quantum walks, see [9, Section 2.4]. A quantum walk is described by a unitary operator on the Hilbert space $\mathcal{H}=\ell^{2}(\mathbb{Z})\otimes\mathbb{C}^{2}$, which models a state space in which a wave packet comes equipped with a “spin” at each integer site. Notice that the elementary tensors of the form $\delta_{n}\otimes e_{\uparrow}$ and $\delta_{n}\otimes e_{\downarrow}$ with $n\in\mathbb{Z}$ comprise an orthonormal basis of $\mathcal{H}$. Here, $\{e_{\uparrow},e_{\downarrow}\}$ denotes the canonical basis of $\mathbb{C}^{2}$. A time-homogeneous quantum walk scenario is given as soon as coins
$C_{n}=\left[\begin{array}[]{cc}c_{n}^{11}&c_{n}^{12}\\ c_{n}^{21}&c_{n}^{22}\end{array}\right]\in\mathbb{U}(2),\quad n\in\mathbb{Z}$ (5.1)
are specified. Assume that $c_{n}^{11},c_{n}^{22}\neq 0$. As one passes from time $t$ to $t+1$, the update rule of the quantum walk is
@@ 5 Quantum walks
$U=\left[\begin{array}[]{ccccccccc}\ddots&\ddots&\ddots&\ddots&&&&&\\ &0&0&c_{0}^{21}&c_{0}^{22}&&&&\\ &c_{-1}^{11}&c_{-1}^{12}&0&0&&&&\\ &&&0&0&c_{1}^{21}&c_{1}^{22}&&\\ &&&c_{0}^{11}&c_{0}^{12}&0&0&&\\ &&&&&0&0&c_{2}^{21}&c_{2}^{22}\\ &&&&&c_{1}^{11}&c_{1}^{12}&0&0\\ &&&&&\ddots&\ddots&\ddots&\ddots\end{array}\right].$ (5.2)
@@
$\mathcal{E}=\left[\begin{array}[]{ccccccccc}\ddots&\ddots&\ddots&\ddots&&&&&\\ &0&0&\bar{\alpha}_{1}&\rho_{1}&&&&\\ &\rho_{-1}&-\alpha_{-1}&0&0&&&&\\ &&&0&0&\bar{\alpha}_{3}&\rho_{3}&&\\ &&&\rho_{1}&-\alpha_{1}&0&0&&\\ &&&&&0&0&\bar{\alpha}_{5}&\rho_{5}\\ &&&&&\rho_{3}&-\alpha_{3}&0&0\\ &&&&&\ddots&\ddots&\ddots&\ddots\end{array}\right].$ (5.3)
The matrix in (5.3) strongly resembles the matrix representation of $U$ in (5.2). However, $\rho_{n}>0$ for all $n$, then (5.2) and (5.3) may not match exactly when $c_{n}^{kk}$ is not real and positive. Indeed, this can be easily resolved by conjugation with a suitable diagonal unitary, as shown in [CMV-quantum-walks].
@@
Let $\Lambda=\mathrm{diag}(\cdots,\lambda_{-1},\lambda_{0},\lambda_{1},\cdots)$. Then one can obtain that
^$\hat{\mathcal{E}}=\Lambda^{*}U\Lambda,$
@@
Notice that the hypotheses $c_{n}^{11},c_{n}^{22}\neq 0$ imply $\rho_{n}>0$ for all $n\in\mathbb{Z}$. According to the analysis above, there is a close relation between a quantum walk and an extended CMV matrix whose Verblunsky coefficients with even index are zero.

### v3

@@ 2.3 Estimate for the Lyapunov Exponent
Assume that for all $\omega\in\mathbb{T}^{d}(p,q)$ and $z\in\mathcal{I_{\alpha}}\subset\mathbb{D}$, $L(\omega,z)>\gamma>0$. Then for any $n\geq 2$,
@@
Obviously, $0\leq L_{n}(\omega,z)\leq C(z)$ for all $n$. Let $k$ be a positive integer such that $k\gamma>16C(z)$. Given $n_{0}>10$, let $l_{0}=[C_{1}(\log n_{0})^{1/\sigma}]$ with some large $C_{1}$. Consider the integers $l_{0},2l_{0},\ldots,2^{k}l_{0}$. Then there exists some $0\leq j<k$ such that (with $l_{j}=2^{j}l_{0}$)
@@
It is not difficult to check that $M(\omega,z;x)$ is conjugate to an $\mathbb{SL}(2,\mathbb{R})$ matrix $T(\omega,z;x)$ through the matrix
$Q=-\frac{1}{1+i}\left(\begin{array}[]{cc}1&-i\\ 1&i\end{array}\right)\in\mathbb{U}(2);$
@@
$T(\omega,z;x)=Q^{*}M(\omega,z;x)Q\in\mathbb{SL}(2,\mathbb{R}),$
@@
Now we apply AP to the matrices $B_{j}=T_{l}(\omega,z;x+(j-1)l\omega)$ for $1\leq j\leq m=[n/l]$ with $\mu=\exp(l\gamma/2)$. Note that $\mu>n^{2}$ if $C_{1}>4/\gamma$. According to LDT, the phases $x$ such that
@@
Based on AP, we have that
@@
We can finally prove this lemma by summing over $2^{k}n$ . ∎
@@ 2.4 Green’s Function and Poisson’s Formula
$\Theta_{n}=\left(\begin{array}[]{cc}\overline{\alpha}_{n}&\rho_{n}\\ \rho_{n}&-\alpha_{n}\end{array}\right).$
@@
$\mathcal{L}_{+}=\left(\begin{matrix}\Theta_{0}&~&~\\ ~&\Theta_{2}&~\\ ~&~&\ddots\end{matrix}\right),\quad\mathcal{M}_{+}=\left(\begin{matrix}\mathbf{1}&~&~\\ ~&\Theta_{1}&~\\ ~&~&\ddots\end{matrix}\right).$
@@
We let $\mathcal{E}_{[a,b]}$ denote the restriction of an extended CMV matrix to a finite interval $a,b]\subset\mathbb{Z}$, defined by
@@
For $z\in\mathbb{C}$, $\beta,\eta\in\partial\mathbb{D}$, we can define the characteristic determinant of matrix $\mathcal{E}^{\beta,\eta}_{[a,b]}$,
@@
Since the equation $\tilde{\mathcal{E}}u=zu$ is equivalent to $(z\mathcal{L}^{*}-\mathcal{M})u=0$, the associated finite-volume Green’s functions are as follows:
@@
According to [22, Proposition 3.8] and [33, Section B.1], for $\beta,\eta\in\partial\mathbb{D}$, the Green’s function has the expression:
@@
From [22, Lemma 3.9], if $u$ satisfies $\tilde{\mathcal{E}}u=zu$, Poisson’s formula reads
@@
Without loss of generality, we restrict the extended CMV matrix to the interval $0,n-1]\subset\mathbb{Z}$. Let $\varphi^{\beta,\eta}_{[0,n-1]}(\omega,z;x)=\det\big{(}z-\mathcal{E}^{\beta,\eta}_{[0,n-1]}\big{)}$ be the characteristic determinant of matrix $\mathcal{E}^{\beta,\eta}_{[0,n-1]}$. For $z\in\partial\mathbb{D}$, it follows from [28, Theorem 2] that the relation between the $n$-step transfer matrix and the characteristic determinant is
@@
Based upon the above analysis, we can get the Green’s function estimate which will be of significance in the proof of AL.
@@
where $C_{0}$ and $\sigma$ are as in LDT. Furthermore, for any $x$ outside $\Omega$,
@@
Without loss of generality, we assume that $0\leq j\leq k\leq n-1$ and $n\gg N_{0}(\gamma)$. Then one can obtain that
@@ 2.5 Semialgebraic Sets
3 Proof of Theorem 1.1
The proof employs a multi-scale analysis based on Green’s function estimates and double resonance exclusion.
The core strategy is to show that for sufficiently large scale $n_{0}$, the finite-volume Green’s function $G_{[0,n_{0}-1]}^{\beta,\eta}$ exhibits exponential off-diagonal decay for most phases $x$. This decay property is the key indicator of localization and is formalized as:
@@
for all $0\leq m_{1},m_{2}\leq n_{0}-1$. To operationalize this condition, we replace (3.1) with an equivalent determinant-based criterion that is more suitable for semi-algebraic analysis:
$\mathop{\sum}\limits_{0\leq m_{1},m_{2}\leq n_{0}-1}$ $e^{2\gamma|m_{1}-m_{2}|}\big{|}\mathrm{det}[(m_{1},m_{2})-\mathrm{minor}\,\,\mathrm{of}\,\,(z-\mathcal{E}_{[0,n_{0}-1]}^{\beta,\eta}(x,\omega))\big{|}^{2}$
@@
This reformulation is crucial as it expresses the Green’s function decay in terms of polynomial conditions on the matrix entries, enabling the application of semi-algebraic geometry tools.
The remainder of the proof will establish that condition (3.2) holds for most phases $x$ and frequencies $\omega$, outside a set of exponentially small measure.
Fix $n_{0}$ and consider the property (3.2). The left-hand side of (3.2) is a polynomial in $(\operatorname{Re}\alpha(x),\operatorname{Im}\alpha(x),\operatorname{Re}z,\operatorname{Im}z)$. Since ${\mathrm{Re}\,}\alpha(x)$ and ${\mathrm{Im}\,}\alpha(x)$ are real analytic functions, we can replace them by the trigonometric polynomials
@@
Fix $\omega\in\mathbb{T}^{d}(p,q)$ and $z\in\mathcal{I_{\alpha}}$. According to Lemma 2.5, the exceptional set $\Omega$ does not only satisfy
@@
Fix $x_{0}\in\mathbb{T}^{d}$. Now we consider the orbit $\{x_{0}+j\omega:|j|\leq n_{1}\}$. In this case, we let $n_{1}=n_{0}^{C}$, where $C$ is a sufficiently large constant.
Apply Lemma 2.7 with $\mathcal{S}=\Omega$, $B=C_{2}n_{0}^{3}$, $\eta=\exp(-C_{0}n^{\sigma})$ and $n=n_{1}$. Then the statement (2.15) implies that except for at most $n_{1}^{1-\delta}$ values of $|j|<n_{1}$, taking $x=x_{0}+j\omega$, one of the intervals $\Lambda$ from (3.6) satisfies (3.7).
Assume that $u=\{u(n)\}_{n\in\mathbb{Z}}$ satisfies the equation
$\tilde{\mathcal{E}}(x_{0},\omega)u=zu$
with $u(0)=1$ and $|u(n)|\lesssim n^{C}$, i.e., $z$ and $u$ are a generalized eigenvalue and a generalized eigenfunction of $\tilde{\mathcal{E}}(x_{0},\omega)$, respectively.
@@
If $j_{0}+1$, $j_{0}+2$, $j_{0}-2$, $j_{0}-1$ satisfy (3.8), one can obtain that
@@
Therefore, fixing $z\in\mathcal{I_{\alpha}}$, if there is a state $u$ with $u(0)=1$, then for any large $n_{0}$, there exists some $j_{0}$, $|j_{0}|<n_{1}=n_{0}^{C}$ for which (3.10) holds.
Denote $\Sigma_{\omega}=\mathop{\cup}\limits_{|j|\leq n_{1}}\sigma(\mathcal{E}_{[-j,j]}^{\beta,\eta}(x_{0}))$. It follows from (3.7) and (3.10) that if
@@
Now consider the interval $\frac{n_{2}}{2},2n_{2}]$ with $n_{2}=n_{0}^{C^{\prime}}$ ($C^{\prime}$ a sufficiently large constant). Suppose that we ensured that
@@
By the paving property from [2, Page 839] or [29, Appendix 7.2], one can obtain that
@@
Now we verify the assumption (3.13).
For $|j|\leq n_{1}$, we consider the set
$\mathcal{B}=\{(\omega,z^{\prime},x):\omega\in\mathbb{T}^{d}(p,q),z^{\prime}\in\sigma(\mathcal{E}_{[-j,j]}^{\beta,\eta}(x_{0})),x\in\Omega(z^{\prime})\}\subset\mathbb{T}^{d}\times\partial\mathbb{D}\times\mathbb{T}^{d}.$
@@
$0,1]^{d}=\mathop{\bigcup}\limits_{l}\Big{(}x_{l}+[0,\frac{1}{n_{2}}]^{d}\Big{)},\,\,l\leq n_{2}^{d}.$
@@
Summing these sets over $j$, $j\leq n_{1}$, we get an $\omega$-set of measure at most $n_{1}n_{2}^{-\frac{1}{12}}<n_{2}^{-\frac{1}{13}}$. This implies that (3.13) holds for $\omega$ outside the $\omega$-set we obtained.
Since we initially fixed $n_{0}$ and set $n_{1}=n_{0}^{C}$, $n_{2}=n_{0}^{C^{\prime}}$, we can denote the above $\omega$-set to be $\Omega_{n_{0}}$. Then we have that
@@
Let $\Omega_{*}=\mathop{\cup}\limits_{n}\mathop{\cap}\limits_{n_{0}>n}\Omega_{n_{0}}$. It follows that $\mathrm{mes}\Omega_{*}=0$ and AL holds for $\omega\in\mathbb{T}^{d}(p,q)\backslash\Omega_{*}$.
@@
4 Application to Quantum Walks
4.1 Quantum Walks and CMV Matrices
Quantum walks share a structural similarity with CMV matrices via unitary equivalence. Now we recall some basic knowledge about quantum walks, see [13, Section 2.4].
A quantum walk is described by a unitary operator on the Hilbert space $\mathcal{H}=\ell^{2}(\mathbb{Z})\otimes\mathbb{C}^{2}$, which models a state space in which a wave packet comes equipped with a “spin” at each integer site. Notice that the elementary tensors of the form $\delta_{n}\otimes e_{\uparrow}$ and $\delta_{n}\otimes e_{\downarrow}$ with $n\in\mathbb{Z}$ comprise an orthonormal basis of $\mathcal{H}$. Here, $\{e_{\uparrow},e_{\downarrow}\}$ denotes the canonical basis of $\mathbb{C}^{2}$. A time-homogeneous quantum walk scenario is given as soon as coins
$C_{n}=\left(\begin{array}[]{cc}c_{n}^{11}&c_{n}^{12}\\ c_{n}^{21}&c_{n}^{22}\end{array}\right)\in\mathbb{U}(2),\quad n\in\mathbb{Z}$ (4.1)
are specified. Assume that $c_{n}^{11},c_{n}^{22}\neq 0$ and $\mathrm{det}\,C_{n}=1$. As one passes from time $t$ to $t+1$, the update rule of the quantum walk is
@@ 4.1 Quantum Walks and CMV Matrices
$U=\left(\begin{array}[]{ccccccccc}\ddots&\ddots&\ddots&\ddots&&&&&\\ &0&0&c_{0}^{21}&c_{0}^{22}&&&&\\ &c_{-1}^{11}&c_{-1}^{12}&0&0&&&&\\ &&&0&0&c_{1}^{21}&c_{1}^{22}&&\\ &&&c_{0}^{11}&c_{0}^{12}&0&0&&\\ &&&&&0&0&c_{2}^{21}&c_{2}^{22}\\ &&&&&c_{1}^{11}&c_{1}^{12}&0&0\\ &&&&&\ddots&\ddots&\ddots&\ddots\end{array}\right).$ (4.2)
@@
$\mathcal{E}=\left(\begin{array}[]{ccccccccc}\ddots&\ddots&\ddots&\ddots&&&&&\\ &0&0&\bar{\alpha}_{1}&\rho_{1}&&&&\\ &\rho_{-1}&-\alpha_{-1}&0&0&&&&\\ &&&0&0&\bar{\alpha}_{3}&\rho_{3}&&\\ &&&\rho_{1}&-\alpha_{1}&0&0&&\\ &&&&&0&0&\bar{\alpha}_{5}&\rho_{5}\\ &&&&&\rho_{3}&-\alpha_{3}&0&0\\ &&&&&\ddots&\ddots&\ddots&\ddots\end{array}\right).$ (4.3)
The matrix in (4.3) strongly resembles the matrix representation of $U$ in (4.2). However, $\rho_{n}>0$ for all $n$, then (4.2) and (4.3) may not match exactly when $c_{n}^{kk}$ is not real and positive. Indeed, this can be easily resolved by conjugation with a suitable diagonal unitary, as shown in [6].
@@
Let $\mathcal{D}=\mathrm{diag}(\cdots,\lambda_{-1},\lambda_{0},\lambda_{1},\cdots)$. Then we can obtain that
$\hat{\mathcal{E}}=\mathcal{D}^{*}U\mathcal{D},$ (4.4)
@@
Notice that the hypotheses $c_{n}^{11},c_{n}^{22}\neq 0$ imply $\rho_{n}>0$ for all $n\in\mathbb{Z}$. According to the analysis above, there is a close relation between a quantum walk and an extended CMV matrix whose Verblunsky coefficients with even index are zero. Moreover, it is worth mentioning that in [8, 10], Cedzich et al. showed that any extended CMV matrix is gauge-equivalent to a quantum walk. [^2] Our gauge equivalence covers general split-step quantum walks as defined in [8, 10].